OpenStudy (lapshaman):

use the given values to find an equation of the form f(x)=ax^2+bx+C f(0)=5, f(1)=10, f(2)=19 I have a test on this tomorrow, so if somebody could explain that would be amazing!

OpenStudy (danjs):

f(0) = 5 ; if x is zero, f(x) = C, so C is 5 ...

OpenStudy (lapshaman):

ok, thanks, but I'm supposed to get a trinomial out of this.....

OpenStudy (danjs):

The other 2 points will give you 2 equations with A and B, you can solve them

OpenStudy (lapshaman):

what do you mean by 2 equations?

OpenStudy (danjs):

f(1) = 10, so ; 10 = a + b + 5 f(2) = 19, so ; 19 = a*2^2 + b*2 + 5

OpenStudy (danjs):

you can figure the 3 missing values out if they give you at least 3 points to make 3 equations with

OpenStudy (lapshaman):

how does that help though, I still have the a and the b...

OpenStudy (danjs):

right, you have 2 equations and 2 unknowns, you can solve that...

OpenStudy (lapshaman):

according to my math book somehow you can by mixing the equations....

OpenStudy (danjs):

simplifying a bit, yeah there are different methods to solve systems like that...a pretty simple way is to solve one for either a or b, then use that expression in the other equation 5 = a + b 14 = 4a + 2b

OpenStudy (danjs):

b = 5 - a from first equation, so 14 = 4a + 2*(5 - a)

OpenStudy (lapshaman):

ok, so we are at 7=2a+b and 5=a+b

OpenStudy (danjs):

yeah, you could reduce it, i forgot, same answer though

OpenStudy (lapshaman):

cool, thanks!!

OpenStudy (danjs):

np, welcome